Two red blood cells each have a mass of 6.35 x 10-14 kg and carry a negative charge spread uniformly over their surfaces The repulsion arising from the excess charge prevents the cells from clumping together. Once cell carries -2.60 pC of charge and the other-2.70 pC, and each cell can be modeled as a sphere 6.40 um in diameter What minimum relative speed v would the red blood cells need when very far away from each other to get close enough to just touch? Ignore viscous drag from the surrounding liquid u= m/s What is the magnitude of the maximum acceleration amas of each cell? m/s amax
Answer #1

the initial speed of blood cells is , v1 = v

the final speed of blood cells is , v2 = 0

initial distance = very far off (infinity)

final speed = 2*radius = 2*(6.40/2)*10-6 = 6.40*10-6 m

The conserving energy is ,

q1*q2/4*pi*\varepsilono*r + 1/2*m*v2 = constant

(q1*q2/4*pi*\varepsilono*\infty) + 2*(1/2*m*v2) = q1*q2/4*pi*\varepsilono*2r + 0

calculate the velocity of the cell ,

m*v2 = q1*q2/4*pi*\varepsilono*2r

v = sqrt[q1*q2/4*pi*\varepsilono*2r*m]

v = sqrt[(-2.60*10-12)*(-2.70*10-12)/4*3.14*8.85*10-12*6.40*10-6*6.35*10-14]

v = 394.20 m/s

These are experiences the maximum acceleration when they just barely touch each other .

so,

m*a = q1*q2/4*pi*\varepsilono*(2r)2

a =  q1*q2/4*pi*\varepsilono*(2r)2*m

a = (-2.60*10-12)*(-2.70*10-12)/4*3.14*8.85*10-12*(6.40*10-6)2*6.35*10-14

a = 2.42*1010 m/s2